Find the slope, if it exists, of the line containing the pair points
(1,6) and (7,-2) The slope m = ------

Answer by checkley77(12569)

(1,6) and (7,-2)
Slope=(y2-y1)/(x2-x1)
Slope=(-2-1)/(7-6)=-3/1=-3 ans. for the slope.